The period of oscillation of an irregularly shaped object is 1.2 s. Its mass is 1.0 kg and it is hanged from one of its ends to oscillate, which is 3.5 cm away from its center of mass. What is its moment of inertia?

1
Expert's answer
2022-01-26T07:41:15-0500
"T=2\pi\sqrt\dfrac I{mrg}→I=\dfrac{mrgT^2}{4\pi^2}=0.0125\text{ kg}·\text m^2."